OpenStudy (anonymous):

Use the Pythagorean Theorem to determine if the three numbers could be the measures of the sides of a right triangle. Assume that the largest is the hypotenuse. 20, 6, 21

OpenStudy (whpalmer4):

Does \(6^2+20^2=21^2\)? If so, the numbers could be the sides of a right triangle. If not, then they couldn't.

OpenStudy (cwrw238):

is 21^2 = 20^2 + 6^2 grab your calculator and try it out